OpenStudy (spinex12):

Expand and evaluate the power. (1.6)2 3.6 3.2 0.8 2.56

OpenStudy (tkhunny):

If it is 1.6^2 = \(1.6^{2}\) Since you have provided no rounding rules, I'm tempted to pick the one with two decimals unless it's WAY our of whack. If it is 1.6*2 = \(1.6\cdot 2\) That's pretty quickly 3.2 What is it and where are you struggling?

OpenStudy (spinex12):

tank you :)

OpenStudy (spinex12):

thank